# Impulsive mixing of stellar populations in dwarf spheroidal galaxies

### v1

@@ 4.3 Impulsive evolution of the density and anisotropy profile
Figure 14: (a) Time evolution of the stellar tracer with $r_{\mathrm{h0}}/r_{\mathrm{mx}}=1/64$, deeply embedded in a cuspy halo. At $t=0$, the stellar tracer is subject to an impulsive perturbation as in Fig. 13. Top panel: (2D) surface density (colour-coded), projected onto the $\{x,y\}$-plane. A black circle with radius $R=r_{\mathrm{h}}$ indicates the (3D) half-light radius. Bottom panel: radial velocity $v_{r}$ and (3D) radius $r$. The number of particles per bin is colour-coded. A cross indicates the location of the half-light radius $r_{\mathrm{h}}$ and the median radial velocity $\langle v_{r}\rangle$. The stellar tracer embedded in a cuspy halo phase mixes rapidly towards an equilibrium configuration. (b) Like (a) but for a stellar tracer deeply embedded in a cored dark matter halo. Phase mixing takes substantially longer than in the cuspy case. (c) Like (a) but for an extended stellar tracer with initial half-light radius $r_{\mathrm{h0}}/r_{\mathrm{mx}}=1/4$, embedded in a cored dark matter halo. In contrast to (b), here, phase mixing progresses rapidly as most stars of the stellar tracer are located outside the density core and hence span a wide range of orbital periods. A video version of this figure is available as an arXiv ancillary file.
@@ 4.5.1 Long-lived oscillatory modes
Some intuition into the cause underlying the slower return to equilibrium of stellar populations deeply embedded in cored subhalos compared to cuspy ones can be gained by focusing on the orbits of individual stars. Strong impulsive perturbations reset the orbital phases of all stars: after decreasing $M/M_{0}$ impulsively, the majority of particles will be located at the pericentre of their new orbit. The return to equilibrium then reduces to phase mixing these new orbits. However, this is less efficient in the centers of cored halos where the range of orbital frequencies is much narrower than in the cuspy case (see the bottom panel of Fig. 4, showing that the crossing times for deeply embedded stellar populations in the harmonic potential of a cored halo all coincide).
@@ 4.5.2 Virial equilibrium
Similar oscillations appear in the evolution of the virial quantities. This is shown in the third row from the top of Fig. 13, where we plot the evolution of the total kinetic energy ($K$, yellow) and the total potential energy ($W$, gray), as well as their sum ($2K+W$, red). Virial equilibrium ($02K+W=$ 0) is reached much more rapidly in the cuspy halo than in the cored one. Consequently, mass modeling based on the assumption of virial equilibrium is prone to substantial systematic error for stellar populations deeply embedded in the harmonic potentials of cored halos. We will illustrate this issue by taking as an example estimators for dynamical masses that make use of the combined measurement of velocity dispersion and size to infer an enclosed dynamical mass (see e.g. Walker et al., 2009; Wolf et al., 2010; Errani et al., 2018). In the fourth panel from the top of Fig. 13, we show the ratio $M_{\mathrm{h}}/M_{\mathrm{est}}$ between (dark matter) mass $M_{\mathrm{h}}=M(<r_{\mathrm{h}})$ enclosed within the half-light radius and a dynamical mass estimate $M_{\mathrm{est}}=r_{\mathrm{h}}\langle\sigma^{2}\rangle/G$. As the stellar population returns to virial equilibrium, the dynamically estimated mass $M_{\mathrm{est}}$ and the actual enclosed mass $M_{\mathrm{h}}$ can differ by up to an order of magnitude.
@@ 4.5.4 Phase space
The phase mixing after an impulsive perturbation leaves clear signatures in phase space (see, e.g. Burger & Zavala 2019 in the context of dark matter core formation, and Laporte et al. 2019 for perturbations to the Galactic disc). For selected simulation snapshots, in Fig. 14 we show the surface brightness $\Sigma_{\star}(x,y)$ (top panels), as well as a projection of phase space using spherical radius $r$ and radial velocity $v_{r}=\bm{v}\cdot\bm{r}/r$ as coordinates (bottom panels). Fig. 14a and Fig. 14b show a stellar population with $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/64$, deeply embedded in a cuspy and a cored halo, respectively. These are the same simulations as shown in Fig. 13 (left panels show the cuspy model, right panels the cored one). The stellar populations return to equilibrium after an impulsive perturbation through inward and outward moving shells of stars. Phase mixing washes out the coherent motions visible in phase space during the first dynamical times after the perturbation. For the case of a stellar population deeply embedded in a cored halo (Figure 14b), the phase space retains structure for many more dynamical times than in the case of the cuspy halo (Figure 14a). Note the caustic coinciding with the initial stellar truncation radius $r_{\mathrm{t0}}$ (separating families of orbits with apocentres above or below $r_{\mathrm{t0}}$) which gives rise to the slight over-density discussed in Sec. 4.3. Finally, Fig. 14c shows the return to equilibrium for a more extended stellar tracer embedded in a cored halo ($r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/4$). Here, phase mixing progresses rapidly, as outside the harmonic potential of the density core, the range of orbital frequencies is much wider than inside the core.
@@ 4.6 Response to linear perturbations
Finally, to make a connection with the results discussed in Rozier & Errani (2024), we turn our attention to the response of a stellar system to impulsive perturbations in the linear regime. As shown in Sec. 3.1, stellar systems subject to adiabatic changes of the underlying potential evolve along curves of $r_{\mathrm{h}}\langle\sigma^{2}\rangle^{1/2}=\text{const}$ towards an equilibrium state. Crucially, the final equilibrium state does not depend on how it was reached but only on the total change of the underlying potential. Whether this change occurred all at once or in several discrete steps does not affect the remnant state, as long as each perturbation occurred on a time scale sufficiently long to be perceived as adiabatic. As a consequence, there is a single evolutionary track for the adiabatic case.
@@
Clearly this does not hold for impulsive perturbations. In Fig. 15, we show the results of an experiment in which we impulsively decrease the mass sourcing the underlying potential to a final value of $M/M_{0}=10^{-1.4}$. We decrease the mass in $n=1$, $2$, $\dots$, $16$ logarithmically spaced steps, and take note of the intermediate equilibrium half-light radii $r_{\mathrm{h}}$ and velocity dispersions $\langle\sigma^{2}\rangle^{1/2}$ of an embedded stellar component. The final state for a single impulsive step ($n=1$) is shown in red, whereas the intermediate and final equilibrium states for $n=2$, $4$, $8$ and $16$ are shown in blue, green, purple, and orange. A dashed gray curve shows the full time evolution towards the final and the intermediate states for the cases with $n=1$ and $n=2$.
@@
We observe that as the number of steps $n$ increases, the final and intermediate states move closer and closer to the adiabatic solution. This behavior is expected; for small changes to the underlying potential, the response of the stellar system is linear, and in the linear case the final state is independent of whether the perturbation is perceived as adiabatic or impulsive.
@@
Figure 15: Equilibrium configurations of half-light radius $r_{\mathrm{h}}$ and average velocity dispersion $\langle\sigma^{2}\rangle^{1/2}$ reached after an impulsive change to the underlying potential. The mass sourcing the potential is decreased to a value of $M/M_{0}=10^{-1.4}$ in $n=1$, $2$, $4$, $8$ and $16$ discrete steps (shown in red, blue, green purple and orange, respectively). The smaller the relative change in potential between two subsequent steps (i.e., the more linear are perturbation and response), the closer is the overall evolution to the adiabatic case. As an example, for the cases of $n=1$ and $n=2$, the full time evolution towards the equilibrium states is depicted using dotted curves.
@@ 5 Summary
    During the adiabatic evolution, each mono-energetic stellar tracer retains centrally isotropic kinematics. Stellar tracers which expand into a region of steepening power-law slope of the underlying (dark matter) profile develop tangentially biased kinematics in their outskirts, and their binding energy distributions slightly widen.
@@
    For the case of power law potentials, stellar binding energy distributions that were initially separated will remain separated during the adiabatic evolution: stellar populations do not mix energetically for adiabatic changes to the underlying potential.
@@
    Consistent with previous work, we show that as impulsive perturbations decrease in amplitude (i.e., as they become more linear), the equilibrium state reached by the stellar population approaches the adiabatic limit.
@@ Acknowledgements
Figure 16: Like Fig. 9, but showing the equilibrium half-light radius $r_{\mathrm{h}}$ and velocity dispersion $\langle\sigma^{2}\rangle^{1/2}$ reached after an impulsive deepening of the underlying potential. The (dark matter) mass sourcing the underlying potential is increased impulsively, in a single step, by $+0.2\,\mathrm{dex}$, $+0.4\,\mathrm{dex}$, $\dots$, $+2\,\mathrm{dex}$. In response to the impulsive increase in mass, the half-light radii contract, and the velocity dispersion increases. The same systematics are observed for stellar tracers embedded in a cored dark matter halo (not shown).
@@
Figure 18: Like Fig. 11, but with energies expressed with respect to the evolved potential minimum. The evolved energy distributions are shown for impulsive reductions of the (dark matter) mass that sources the potential (top three panels), the initial conditions (central panel), and for impulsive deepening of the underlying potential (bottom three panels, see Appendix A for a discussion of the equilibrium structural parameters for this scenario). For both the case of impulsive (dark matter) mass increase and decrease, the stellar binding energy distributions widen with respect to the initial conditions, and may overlap, mixing the initially energetically separated populations.
Appendix A Impulsive deepening of the underlying potential
In this appendix, we discuss the response of spherical stellar populations to an impulsive deepening of the underlying potential. We consider stellar populations that in the initial conditions are mono-energetic and have isotropic kinematics, with initial half-light radii of $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1$, $1/4$, $1/16$ and $1/64$. The numerical setup is identical to that used in Sec. 4. All simulations start with the same initial condition. We then impulsively increase the mass $M$ which sources the potential, in a single step, by $+0.2\,\mathrm{dex}$, $+0.4\,\mathrm{dex}$, $\dots$, $+2.0\,\mathrm{dex}$, respectively. As an example, here we discuss the evolution of stellar populations in cuspy dark matter halos, but the same systematics hold also for the case of stellar tracers embedded in cored halos.
@@ Appendix A Impulsive deepening of the underlying potential
Figure 19: Number density profiles $r^{2}\nu_{\star}(r)$ for mono-energetic stellar tracers with purely radial orbits (Eq. B2), embedded in a cuspy (solid curve, Eq. 4) and a cored (dashed curve, Eq. 21) dark matter halo. Profiles are shown for four different values of the stellar truncation radius $r_{\mathrm{t}}/r_{\mathrm{mx}}=1/10$, $1$, $10$, $100$. The profiles asymptotically behave like $\nu_{\star}\propto r^{-2}$ for $r\rightarrow 0$, and have a caustic at the radius $r_{\mathrm{t}}$ caused by the pileup of stars at the apocentre $r_{\mathrm{t}}$.
@@ Appendix B Mono-energetic tracer with purely radial orbits
where by $\mathcal{E}$ we denote the energy (Eq. 2) and by $J$ the magnitude of the angular momentum (Eq. 27). The distribution function above is identical to zero, unless a star has angular momentum zero and energy $\mathcal{E}_{\star}$. By integrating the distribution function over all velocities we find the density profile given by Eq. B2. One may reach the same result by considering that the number of stars $\mathrm{d}N_{\star}$ located at a radius $r$ is proportional to the time each star spends at that radius along its orbit, $\mathrm{d}t=\mathrm{d}r/v(r)$, where $v(r)$ is the star’s velocity. Hence $\mathrm{d}N_{\star}=4\pi r^{2}\nu_{\star}(r)\mathrm{d}r\propto\mathrm{d}r/v(r)$, and the stellar number density profile reads
$\nu_{\star}(r)=\nu_{\mathrm{s}}~{}~{}(r/a)^{-2}~{}\left[\mathcal{E}_{\star}-\psi(r)\right]^{-1/2}~{},$ (B2)
where by $\psi(r)$ we denote the underlying gravitational potential (Eq. 3), $a$ is a scale radius (chosen here to correspond to the scale radius of the underlying dark matter halo), and $\nu_{\mathrm{s}}$ is a stellar scale density. This mono-energetic density profile is cuspy with $\mathrm{d}\ln\nu_{\star}(r)/\mathrm{d}\ln r\rightarrow-2$ for $r\rightarrow 0$, and has a caustic at the radius $r_{\mathrm{t}}$ where $\psi(r_{\mathrm{t}})=\mathcal{E}_{\star}$. This radius is identical to the truncation radius discussed in Sec. 2 (Eq. 18, 23) beyond which the stellar density is zero. The caustic is caused by the (radial) velocity of stars dropping to zero at their apocentre $r_{\mathrm{t}}$. Fig. 19 shows four example stellar profiles following Eq. B2 with $r_{\mathrm{t}}/r_{\mathrm{mx}}=1/10$, $1$, $10$, $100$, where by $r_{\mathrm{mx}}$ we denote the radius of peak circular velocity of the underlying dark matter halo. Solid and dashed curves correspond to stellar profiles embedded in a cuspy (Eq. 4) and a cored (Eq. 21) dark matter halo, respectively.
We compute the scale density $\nu_{s}$ by requiring that $N_{\star}=4\pi\int_{0}^{r_{\mathrm{t}}}r^{2}\nu_{\star}(r)\mathrm{d}r$. For a stellar tracer of energy $\mathcal{E}_{\star}$ embedded in a cuspy ($\gamma=1$) Hernquist halo with scale radius $a=r_{\mathrm{mx}}$, we find
@@
which for the case of a stellar component embedded deeply in the constant-density core, $\mathcal{E}_{\star}\ll 1$, reduces to $\nu_{s}\approx N_{\star}/\left(2\pi^{2}a^{3}\right)$.
In Sec. 4.3, we use a superposition of stellar profiles following Eq. B2 to model the outer density profiles of stellar systems that were affected by impulsive perturbations.

### v2

@@ 4.3 Impulsive evolution of the density and anisotropy profile
Figure 14: (a) Time evolution of the stellar tracer with $r_{\mathrm{h0}}/r_{\mathrm{mx}}=1/64$, deeply embedded in a cuspy halo. At $t=0$, the stellar tracer is subject to an impulsive perturbation as in Fig. 13. Top panel: (2D) surface density (colour-coded), projected onto the $\{x,y\}$-plane. A black circle with radius $R=r_{\mathrm{h}}$ indicates the (3D) half-light radius. Bottom panel: radial velocity $v_{r}$ and (3D) radius $r$. The number of particles per bin is colour-coded. A cross indicates the location of the half-light radius $r_{\mathrm{h}}$ and the median radial velocity $\langle v_{r}\rangle$. The stellar tracer embedded in a cuspy halo phase mixes rapidly toward an equilibrium configuration. (b) Like (a) but for a stellar tracer deeply embedded in a cored dark matter halo. Phase mixing takes substantially longer than in the cuspy case. (c) Like (a) but for an extended stellar tracer with initial half-light radius $r_{\mathrm{h0}}/r_{\mathrm{mx}}=1/4$, embedded in a cored dark matter halo. In contrast to (b), here, phase mixing progresses rapidly as most stars of the stellar tracer are located outside the density core and hence span a wide range of orbital periods. For videos supplementing this figure see Appendix C.
@@ 4.5.1 Long-lived oscillatory modes
Some intuition into the cause underlying the slower return to equilibrium of stellar populations deeply embedded in cored subhalos compared to cuspy ones can be gained by focusing on the orbits of individual stars. Strong impulsive perturbations reset the orbital phases of all stars: after decreasing $M/M_{0}$ impulsively, the majority of particles will be located at the pericentre of their new orbit. The return to equilibrium of the stellar tracer particles then reduces to phase mixing these new orbits. This process is less efficient in the central regions of cored halos, where the range of orbital frequencies is much narrower than in cuspy ones. This is illustrated in the bottom panel of Fig. 4, which shows that the crossing times for deeply embedded stellar populations in the harmonic potential of a cored halo are nearly identical (see also Banik & van den Bosch 2021b for a discussion of phase mixing in cored halos, and Sridhar 1989 for oscillatory modes in stellar systems).
Note that in the present work, we describe the dynamical evolution of mass-less tracer particles in a background potential, and phase mixing is the sole process that allows the system of tracer particles to return to an equilibrium state. In contrast, self-gravitating systems can exhibit mode damping that is facilitated by energy exchange between particles (see e.g. Weinberg 1994b, Kandrup (1998), Nelson & Tremaine 1999, Fouvry & Prunet 2022 and more recently Petersen et al. 2024, Dattathri et al. 2025 for discussions of Landau damping in self-gravitating stellar systems).
@@ 4.5.2 Virial equilibrium
Oscillations also appear in the evolution of the virial quantities. This is shown in the third row from the top of Fig. 13, where we plot the evolution of the total kinetic energy ($K$, yellow) and the total potential energy ($W$, gray), as well as their sum ($2K+W$, red). Virial equilibrium ($2K+W=0$) is reached much more rapidly in the cuspy halo than in the cored one. Consequently, mass modeling based on the assumption of virial equilibrium is prone to substantial systematic error for stellar populations deeply embedded in the harmonic potentials of cored halos. We will illustrate this issue by taking as an example estimators for dynamical masses that make use of the combined measurement of velocity dispersion and size to infer an enclosed dynamical mass (see e.g. Walker et al., 2009; Wolf et al., 2010; Errani et al., 2018). In the fourth panel from the top of Fig. 13, we show the ratio $M_{\mathrm{h}}/M_{\mathrm{est}}$ between (dark matter) mass $M_{\mathrm{h}}=M(<r_{\mathrm{h}})$ enclosed within the half-light radius and a dynamical mass estimate $M_{\mathrm{est}}=r_{\mathrm{h}}\langle\sigma^{2}\rangle/G$. As the stellar population returns to virial equilibrium, the dynamically estimated mass $M_{\mathrm{est}}$ and the actual enclosed mass $M_{\mathrm{h}}$ can differ by up to an order of magnitude.
@@ 4.5.4 Phase space
The phase mixing after an impulsive perturbation leaves clear signatures in phase space (see, e.g. Burger & Zavala 2019 in the context of dark matter core formation, and Laporte et al. 2019; Banik et al. 2022, 2023 for perturbations to the Galactic disc). For selected simulation snapshots, in Fig. 14 we show the surface brightness $\Sigma_{\star}(x,y)$ (top panels), as well as a projection of phase space using spherical radius $r$ and radial velocity $v_{r}=\bm{v}\cdot\bm{r}/r$ as coordinates (bottom panels). Fig. 14a and Fig. 14b show a stellar population with $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/64$, deeply embedded in a cuspy and a cored halo, respectively. These are the same simulations as shown in Fig. 13 (left panels show the cuspy model, right panels the cored one). The stellar populations return to equilibrium after an impulsive perturbation through inward and outward moving shells of stars. Phase mixing washes out the coherent motions visible in phase space during the first dynamical times after the perturbation. For the case of a stellar population deeply embedded in a cored halo (Figure 14b), the phase space retains structure for many more dynamical times than in the case of the cuspy halo (Figure 14a). Note the caustic coinciding with the initial stellar truncation radius $r_{\mathrm{t0}}$ (separating families of orbits with apocentres above or below $r_{\mathrm{t0}}$) which gives rise to the slight overdensity discussed in Sec. 4.3. Finally, Fig. 14c shows the return to equilibrium for a more extended stellar tracer embedded in a cored halo ($r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1/4$). Here, phase mixing progresses rapidly, as outside the harmonic potential of the density core, the range of orbital frequencies is much wider than inside the core.
@@ 4.6 Response to linear perturbations
Finally, to make a connection with the results discussed in Rozier & Errani (2024), we turn our attention to the response of a stellar system to impulsive perturbations in the linear regime. As shown in Sec. 3.1, stellar systems subject to adiabatic changes of the underlying potential evolve along curves of $r_{\mathrm{h}}\langle\sigma^{2}\rangle^{1/2}=\text{const}$ toward an equilibrium state. Crucially, the final equilibrium state does not depend on how it was reached but only on the total change of the underlying potential. Whether this change occurred all at once or in several discrete steps does not affect the remnant state, as long as each perturbation occurred on a time scale sufficiently long to be perceived as adiabatic. As a consequence, there is a single evolutionary track for the adiabatic case.
@@
Clearly this does not hold for impulsive perturbations. In Fig. 15, we show the results of an experiment in which we impulsively decrease the mass sourcing the underlying potential to a final value of $M/M_{0}=10^{-1.4}$. We decrease the mass in $n=1$, 22, $\dots$, $16$ logarithmically spaced steps, and take note of the intermediate equilibrium half-light radii $r_{\mathrm{h}}$ and velocity dispersions $\langle\sigma^{2}\rangle^{1/2}$ of an embedded stellar component. The final state for a single impulsive step ($n=1$) is shown in red, whereas the intermediate and final equilibrium states for $n=2$, 44, 88 and $16$ are shown in blue, green, purple, and orange. A dashed gray curve shows the full time evolution toward the final and the intermediate states for the cases with $n=1$ and $n=2$.
@@
We observe that as the number of steps $n$ increases, the final and intermediate states move closer and closer to the adiabatic solution. This behavior is consistent with the analytical results discussed in Rozier & Errani (2024): for stellar systems that are only slightly out of equilibrium, with both the perturbation and the response in the linear regime and non-resonant, when neglecting second-order terms, Rozier & Errani (2024) found that the system’s global properties in the final equilibrium state are independent of the detailed time evolution of the perturbation. In other words, in the linear, non-resonant regime, the global properties of the final equilibrium state reached by a stellar system in response to a perturbation are independent of whether the perturbation is perceived as adiabatic or impulsive. In the numerical experiments presented in Fig. 15, for small consecutive impulsive changes to the underlying potential, both the perturbation to the stellar component as well as its response are approximately linear, and we recover an evolution that coincides with the adiabatic case.
@@
Figure 15: Equilibrium configurations of half-light radius $r_{\mathrm{h}}$ and average velocity dispersion $\langle\sigma^{2}\rangle^{1/2}$ reached after an impulsive change to the underlying potential. The mass sourcing the potential is decreased to a value of $M/M_{0}=10^{-1.4}$ in $n=1$, 22, 44, 88 and 1616 discrete steps (shown in red, blue, green, purple and orange, respectively). The smaller the relative change in potential between two subsequent steps (i.e., the more linear are perturbation and response), the closer is the overall evolution is to the adiabatic case. As an example, for the cases of $n=1$ and $n=2$, the full time evolution toward the equilibrium states is depicted using dotted curves.
@@ 5 Summary
    During the adiabatic evolution, each mono-energetic stellar tracer retains centrally isotropic kinematics. Stellar tracers that expand into a region of steepening power-law slope of the underlying (dark matter) profile develop tangentially biased kinematics in their outskirts, and their binding energy distributions slightly widen.
@@
    For the case of power-law potentials, stellar binding energy distributions that were initially separated will remain separated during the adiabatic evolution: stellar populations do not mix energetically for adiabatic changes to the underlying potential.
@@
    Consistent with previous work, we show that as impulsive perturbations decrease in amplitude (i.e., as they become more linear), the equilibrium state reached by the stellar population approaches that of the adiabatic case.
@@ Acknowledgements
Appendix A Impulsive deepening of the underlying potential
In this appendix, we discuss the response of spherical stellar populations to an impulsive deepening of the underlying potential. We consider stellar populations that in the initial conditions are mono-energetic and have isotropic kinematics, with initial half-light radii of $r_{\mathrm{h0}}/r_{\mathrm{mx0}}=1$, $1/4$, $1/16$ and $1/64$. The numerical setup is identical to that used in Sec. 4. All simulations start with the same initial condition. We then impulsively increase the mass $M$ which sources the potential, in a single step, by $0.2\,\mathrm{dex}$, $0.4\,\mathrm{dex}$, $\dots$, $2.0\,\mathrm{dex}$, respectively. As an example, here we discuss the evolution of stellar populations in cuspy dark matter halos, but the same systematics hold also for the case of stellar tracers embedded in cored halos.
Figure 16: Like Fig. 9, but showing the equilibrium half-light radius $r_{\mathrm{h}}$ and velocity dispersion $\langle\sigma^{2}\rangle^{1/2}$ reached after an impulsive deepening of the underlying potential. The (dark matter) mass sourcing the underlying potential is increased impulsively, in a single step, by $0.2\,\mathrm{dex}$, $0.4\,\mathrm{dex}$, $\dots$, $2\,\mathrm{dex}$. In response to the impulsive increase in mass, the half-light radii contract, and the velocity dispersion increases. The same systematics are observed for stellar tracers embedded in a cored dark matter halo (not shown).
@@ Appendix A Impulsive deepening of the underlying potential
Figure 18: Like Fig. 11, but with energies expressed with respect to the evolved potential minimum. The evolved energy distributions are shown for impulsive reductions of the (dark matter) mass that sources the potential (top three panels), the initial conditions (central panel), and for impulsive deepening of the underlying potential (bottom three panels; see Appendix A for a discussion of the equilibrium structural parameters for this scenario). For both the cases of impulsive (dark matter) mass increase and decrease, the stellar binding energy distributions widen with respect to the initial conditions, and may overlap, mixing the initially energetically separated populations.
@@ Appendix B Mono-energetic tracer with purely radial orbits
where by $\mathcal{E}$ we denote the energy (Eq. 2) and by $J$ we denote the magnitude of the angular momentum (Eq. 27). The distribution function above is identical to zero, unless a star has angular momentum zero and energy $\mathcal{E}_{\star}$. By integrating the distribution function over all velocities, we find the density profile given by Eq. B2. One may reach the same result by considering that the number of stars $\mathrm{d}N_{\star}$ located at a radius $r$ is proportional to the time each star spends at that radius along its orbit, $\mathrm{d}t=\mathrm{d}r/v(r)$, where $v(r)$ is the star’s velocity. Hence, $\mathrm{d}N_{\star}=4\pi r^{2}\nu_{\star}(r|\mathcal{E}_{\star})\mathrm{d}r\propto\mathrm{d}r/v(r)$, and the stellar number density profile reads
$\nu_{\star}(r|\mathcal{E}_{\star})=\begin{dcases}\nu_{\mathrm{s}}~(a/r)^{2}\big/\sqrt{\mathcal{E}_{\star}-\psi(r)}&\text{if }\psi(r)<\mathcal{E}_{\star}\\ 0~&\text{otherwise,}\end{dcases}$ (B2)
where by $\psi(r)$ we denote the underlying gravitational potential (Eq. 3), $a$ is a scale radius (chosen here to correspond to the scale radius of the underlying dark matter halo), and $\nu_{\mathrm{s}}$ is a stellar scale density. This mono-energetic density profile is cuspy with $\mathrm{d}\ln\nu_{\star}(r|\mathcal{E}_{\star})/\mathrm{d}\ln r\rightarrow-2$ for $r\rightarrow 0$, and has a caustic at the radius $r_{\mathrm{t}}$ where $\psi(r_{\mathrm{t}})=\mathcal{E}_{\star}$. This radius is identical to the truncation radius discussed in Sec. 2 (Eq. 18, 23) beyond which the stellar density is zero. The caustic is caused by the (radial) velocity of stars dropping to zero at their apocentre $r_{\mathrm{t}}$. Fig. 19 shows four example stellar profiles following Eq. B2 with $r_{\mathrm{t}}/r_{\mathrm{mx}}=1/10$, $1$, 1010, 100100, where by $r_{\mathrm{mx}}$ we denote the radius of peak circular velocity of the underlying dark matter halo. Solid and dashed curves correspond to stellar profiles embedded in a cuspy (Eq. 4) and a cored (Eq. 21) dark matter halo, respectively.
Figure 19: Number density profiles $r^{2}\nu_{\star}(r|\mathcal{E}_{\star})$ for mono-energetic stellar tracers with purely radial orbits (Eq. B2), embedded in a cuspy (solid curve, Eq. 4) and a cored (dashed curve, Eq. 21) dark matter halo. Profiles are shown for four different values of the stellar truncation radius $r_{\mathrm{t}}/r_{\mathrm{mx}}=1/10$, $1$, 1010, 100100. The profiles asymptotically behave like $\nu_{\star}\propto r^{-2}$ for $r\rightarrow 0$, and have a caustic at the radius $r_{\mathrm{t}}$ caused by the pileup of stars at the apocentre $r_{\mathrm{t}}$.
Figure 20: Animated figure supplementing the left-hand panel of Fig. 13. The video shows the time evolution of the half-light radius and the velocity dispersion of a stellar tracer population in response to an impulsive perturbation in a cuspy dark matter halo. File size $0.9\,\mathrm{MB}$, dimensions $3840\times 1920\,\mathrm{px}^{2}$, duration $33\,\mathrm{s}$. The animated figure is available in the arXiv ancillary files.
Figure 21: Animated figure supplementing the right-hand panel of Fig. 13. The video shows the time evolution of the half-light radius and the velocity dispersion of a stellar tracer population in response to an impulsive perturbation in a cored dark matter halo. File size $1.3\,\mathrm{MB}$, dimensions $3840\times 1920\,\mathrm{px}^{2}$, duration $33\,\mathrm{s}$. The animated figure is available in the arXiv ancillary files.
We compute the scale density $\nu_{s}$ by requiring that $N_{\star}=4\pi\int_{0}^{r_{\mathrm{t}}}r^{2}\nu_{\star}(r|\mathcal{E}_{\star})\mathrm{d}r$. For a stellar tracer of energy $\mathcal{E}_{\star}$ embedded in a cuspy ($\gamma=1$) Hernquist halo with scale radius $a=r_{\mathrm{mx}}$, we find
@@
which, for the case of a stellar component embedded deeply in the constant-density core, $\mathcal{E}_{\star}\ll 1$, reduces to $\nu_{s}\approx N_{\star}/\left(2\pi^{2}a^{3}\right)$.
In Sec. 4.3, we use a superposition of stellar profiles following Eq. B2 to model the outer density profiles of stellar systems that were affected by impulsive perturbations. The superposition is computed by weighting the mono-energetic profiles $\nu_{\star}(r)$ by differential energy distribution $\mathrm{d}N_{\star}/\mathrm{d}\mathcal{E}$ of Fig. 11, integrating from the most-bound state $\mathcal{E}_{\star}=0$ to the bound/unbound limit $\mathcal{E}_{\star}=1$,
$\nu_{\star}(r)=\int_{0}^{1}\frac{\mathrm{d}N_{\star}}{\mathrm{d}\mathcal{E}_{\star}}~\nu_{\star}(r|\mathcal{E}_{\star})~\mathrm{d}\mathcal{E}_{\star}~,$ (B5)
where each mono-energetic $\nu_{\star}(r|\mathcal{E}_{\star})$ is normalized here so that $1=4\pi\int_{0}^{\infty}r^{2}\nu_{\star}(r|\mathcal{E}_{\star})\mathrm{d}r$.
For stellar energy distributions that extend all the way to the bound/unbound limit, at large radii, the log-slope of this model density profile steepens to a value of $\mathrm{d}\ln\nu_{\star}(r)/\mathrm{d}\ln r\rightarrow-4$. This can be shown by noting that, for $\mathcal{E}_{\star}\sim 1$, the density profile of a mono-energetic stellar tracer scales as $\nu_{\star}(r|\mathcal{E}_{\star})\propto r^{-2}(1-\mathcal{E}_{\star})^{3/2}/\sqrt{\mathcal{E}_{\star}-\psi(r)}$. At each radius $r\gg a$ within the density tail, only stars on orbits with energy $\mathcal{E}_{\star}\geq\psi(r)$ contribute to the local density. For a stellar energy distribution $\mathrm{d}N_{\star}/\mathrm{d}\mathcal{E}_{\star}\approx\text{const}$ that is flat at the bound/unbound limit $\mathcal{E}_{\star}=1$, the superposition of mono-energetic stellar density profiles at $r\gg a$ reads
$\nu_{\star}(r)\!\stackrel{{\scriptstyle r\gg a}}{{\propto}}\!\frac{1}{r^{2}}\!\!\int\limits_{\psi(r)}^{1}\!\!\frac{(1\!-\!\mathcal{E}_{\star})^{3/2}}{\sqrt{\mathcal{E}_{\star}\!-\!\psi(r)}}\,\mathrm{d}\mathcal{E}_{\star}\!\propto\!\frac{[\psi(r)\!-\!1]^{2}}{r^{2}}\!\propto\!\frac{1}{r^{4}}.$ (B6)
Dotted curves in Fig. 12 (Sec. 4.3) show stellar density profiles that are modeled as a superposition of mono-energetic densities (Eq. B2) weighted by the respective stellar energy distribution of Eq. 40 (shown as a dotted curve in Fig. 11). For stellar tracers whose energy distributions extend all the way to the bound/unbound limit, the log-slope $\mathrm{d}\ln\nu_{\star}(r)/\mathrm{d}\ln r$ approaches 4-4 for $r\gg a$, consistent with the analytical estimate of Eq. B6.
Figure 22: Animated figure supplementing Fig. 14a. The video shows the time evolution of stellar tracer particles in configuration space $\{x,y\}$ and in phase space $\{r,v_{r}\}$ in response to an impulsive perturbation in a cuspy dark matter halo. File size $9.2\,\mathrm{MB}$, dimensions $3840\times 1920\,\mathrm{px}^{2}$, duration $29\,\mathrm{s}$. The animated figure is available in the arXiv ancillary files.
Figure 23: Animated figure supplementing Fig. 14b. The video shows the time evolution of stellar tracer particles in configuration space $\{x,y\}$ and in phase space $\{r,v_{r}\}$ in response to an impulsive perturbation in a cored dark matter halo. File size $8.0\,\mathrm{MB}$, dimensions $3840\times 1920\,\mathrm{px}^{2}$, duration $29\,\mathrm{s}$. The animated figure is available in the arXiv ancillary files.
Appendix C Supplementary animated figures
This appendix contains brief descriptions of the animated figures 20, 21, 22 and 23, provided on the journal website.